### v2

# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

Abstract
We build a finite volume scheme for the scalar conservation law $\partial_{t}u+\partial_{x}(H(x,u))=0$ with initial condition $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ for a wide class of flux function $H$, convex with respect to the second variable. The main idea for the construction of the scheme is to use the theory of discontinuous flux. We prove that the resulting approximating sequence converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$, for all $p\in[1,+\infty\mathclose{[}$, to the entropy solution.
11footnotetext: abraham.sylla@unibs.it
Department of Information Engineering, University of Brescia (Italy)
ORCID number: $0$-0003-1784-4878
2020 AMS Subject Classification: 65M08, 65M12, 35L65.
Keywords: Scalar Conservation Law, Compactly Space Dependent Flux, Finite Volume Scheme, Discontinuous Flux, Well-Balanced Scheme.
1 Introduction
Consider the following Cauchy problem for $x$-dependent scalar conservation law:
$\partial_{t}u(t,x)+\partial_{x}\left(H(x,u(t,x))\right)$ $=0\qquad(t,x)\in\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$
$u(0,x)$ $=u_{o}(x)\qquad x\in\mathbb{R}.$ (CL)
Equations of this type often occur, for instance, in the modeling of physical phenomena related to traffic flow [40, 31, 20], porous media [19, 21] or sedimentation problems [13, 14].
It is known that solutions of (1) are discontinuous, regardless of the regularity of the data. They are to be understood in the entropy sense of [27]. The following quantity often recurs below,
$\forall x,u,k\in\mathbb{R},\quad\Phi(x,u,k)\coloneq\mathop{\rm sgn}(u-k)\;(H(x,u)-H(x,k)).$ (1.1)
Definition 1.1.
Let $H\in\mathbf{C}^{1}(\mathbb{R}^{2},\mathbb{R})$ and $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We say that $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ is an entropy solution to (1) if for all $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and $k\in\mathbb{R}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-k|\partial_{t}\varphi+\Phi(x,u,k)\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}-\int_{0}^{+\infty}\int_{\mathbb{R}}\mathop{\rm sgn}(u-k)\partial_{x}H(x,k)\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$+\int_{\mathbb{R}}|u_{o}(x)-k|\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (1.2)
In appearance, Definition 1.1 is weaker than the classical [27, Definition 1] since it does not require the existence of a strong trace at the initial time. It is in particular more manageable to limit process. Nevertheless, it can be shown that if $H\in\mathbf{C}^{3}(\mathbb{R}^{2},\mathbb{R})$, then Definition 1.1 ensures that each entropy solution admits a representative belonging to $\mathbf{C}(\mathbb{R}^{+},\mathbf{L}_{\mathbf{loc}}^{1}(\mathbb{R},\mathbb{R}))$, see [11, Theorem 2.6].
In [27], the author proved existence, uniqueness and stability with the respect to the initial datum for (1) in the framework of entropy solutions under, among others, the growth assumptions
$\partial_{u}H\in\mathbf{L}^{\infty}(\mathbb{R}^{2},\mathbb{R})\quad\text{and}\quad\sup_{(x,u)\in\mathbb{R}^{2}}\left(-\partial_{xu}^{2}H(x,u)\right)<+\infty,$ (1.3)
highlighting that the space dependency in (1) was treated as a source term. In his paper, Kruzhkov proved existence using the vanishing viscosity method, that is, through a parabolic regularization of (1). The following example motivates the necessity to relax (1.3) in the 1D scalar case.
Example 1.1.
Fix positive constants $X,V_{1},V_{2},R_{1},R_{2}$, and let $\mathcal{V},\rho\in\mathbf{C}^{3}(\mathbb{R},\mathopen{]}0,+\infty\mathclose{[})$ be such that $\mathcal{V}(x)=V_{1}$, resp. $\rho(x)=R_{1}$, for $x<-X$ and $\mathcal{V}(x)=V_{2}$, resp. $\rho(x)=R_{2}$, for $x>X$. Define
$H(x,u)\coloneq\mathcal{V}(x)\;u\left(1-\frac{u}{\rho(x)}\right).$
Then with this flow, (1) is the so-called “LWR” (Lightill-Whitham, Richards) model [28, 34] for a flow of vehicles described by their density $u$ along a rectilinear road with maximal speed, resp. density, smoothly varying from $V_{1}$ to $V_{2}$, resp. from $R_{1}$ to $R_{2}$. Clearly, $H$ does not satisfy (1.3).
Remark 1.1.
A priori, Kruzhkov results do not apply in the case of Example 1.1. For completeness, let us however mention that a truncation argument could be used to extend them if the initial datum takes values between the stationary solutions $u(t,x)=0$ and $u(t,x)=\rho(x)$.
Recently, the authors of [11] developed an alternative framework to tackle (1), one not requiring (1.3) and inspired by Example 1.1:
$\mathbf{Smoothness:}$ $\quad H\in\mathbf{C}^{3}(\mathbb{R}^{2},\mathbb{R}).$ (C3)
$\mathbf{Compact\leavevmode\nobreak\ Heterogeneity:}$ $\quad\exists X>0,\;\forall(x,u)\in\mathbb{R}^{2},\quad|x|\geq X\implies\partial_{u}H(x,u)=0.$ (CH)
$\mathbf{Strong\leavevmode\nobreak\ Convexity:}$ $\quad\forall x\in\mathbb{R},\;u\mapsto\partial_{u}H(x,u)\;\text{is an increasing}$ (CVX)
$\mathbf{C}^{1}$-diffeomorphism of $\mathbb{R}$ onto itself.
Assumption (CVX) ensures that for all $x\in\mathbb{R}$, the mapping $u\mapsto H(x,u)$ is strictly convex. Naturally, it can be replaced by the strong concavity assumption: for all $x\in\mathbb{R}$, $u\mapsto\partial_{u}H(x,u)$ is a decreasing $\mathbf{C}^{1}$-diffeomorphism of $\mathbb{R}$ onto itself.
The condition (CH) expresses the compact spatial heterogeneity of $H$ and is not, apparently, common in the context of conservation laws. We expect that it might be relaxed.
Like we previously mentioned, Assumptions (C3)–(CVX) comprise flows that do not fit in the classical Kruzhkov framework and that are relevant from the modeling point of view, for instance, the flow of Example 1.1 (in the concave case). On the other hand, Kruzhkov results apply to general balance laws in several space dimensions. Let us recall the following.
Theorem 1.2.
[11, Theorem 2.6 and Theorem 2.18] Assume that $H$ satisfies (C3)–(CVX). Then for all $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, the Cauchy problem (1) admits a unique entropy solution $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
Moreover, if $v_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and if $v$ is the associated entropy solution, then there exists $L>0$ such that for all $R>0$ and $t>0$,
$\int_{|x|\leq R}|u(t,x)-v(t,x)|\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}|u_{o}(x)-v_{o}(x)|\mathinner{\mathrm{d}{x}}$
$\int_{|x|\leq R}(u(t,x)-v(t,x))^{+}\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}(u_{o}(x)-v_{o}(x))^{+}\mathinner{\mathrm{d}{x}}.$
The proof of existence in [11], like in [27], relies on the vanishing viscosity method, but there, the authors exploit the correspondence with the Hamilton-Jacobi equation (and its parabolic approximation), going back and forth between the two frameworks, and gathering information for both equations at each step.
Remark 1.2.
To be precise, let us mention that in [11], the convexity assumption (CVX) is relaxed to a uniform coercivity assumption coupled to a genuine nonlinearity assumption:
$\mathbf{Uniform\leavevmode\nobreak\ Coercivity:}$ $\quad\forall h\in\mathbb{R},\;\exists\;\mathcal{U}_{h}\in\mathbb{R},\;\forall(x,u)\in\mathbb{R}^{2},$
$\quad|H(x,u)|\leq h\implies|u|\leq\mathcal{U}_{h}.$ (UC)
[…]
has empty interior. (WGNL)
[…]
$\forall x,u\in\mathbb{R},\;H(x,u)\geq\phi(|u|)\quad\text{and}\quad\frac{\phi(r)}{r}{\ \underset{r\to+\infty}{\longrightarrow}\ }+\infty.$
The convexity was used in [12] to characterize, for (1), the attainable set and the set of initial data evolving at a prescribed time into a prescribed profile.
An alternative to the vanishing viscosity method to construct solutions to (1) is to build and prove the convergence of finite volume schemes. The most recent results include the works of [18, 17, 9] (in the case $H(x,u)=\theta(x)f(u)$), [7] (convergence and error analysis), [8] (with a source term) or [39] (on a bounded domain). In all these works, the flux function enters the framework of Kruzhkov by satisfying either (1.3), or the stronger requirement $\partial_{x}H\equiv 0$.
The aim of this paper is to build a finite volume scheme for (1) and prove its convergence to the entropy solution under the assumptions (C3)–(CVX). As a byproduct, we provide an alternate existence result for the Cauchy problem (1), one that does not rely on the vanishing viscosity method. However, let us stress that the main result of the paper is a convergence result and not an existence result. To the author’s knowledge, no convergence result is available for flux functions that do not verify (1.3).
In the convergence analysis of any standard finite volume scheme for (1), a difficulty to tackle is the obtaining of a priori $\mathbf{L}^{\infty}$ bounds for the approximating sequence. Indeed, because of the space dependency, constants are not steady states anymore, therefore the maximum principle breaks down. In [18, 17, 9, 7, 8, 39], this issue was overcome either with the requirement $\partial_{x}H\equiv 0$, ensuring that constants are steady states, or with Assumption (1.3), which provides a time dependent $\mathbf{L}^{\infty}$ bounds of the numerical solution. We choose a new approach, relying on the theory of discontinuous flux, see [1, 2, 3, 4, 13, 19, 24, 26, 30, 37] and the references therein. For the construction of the scheme, the space dependency of $H$ is discretized, and the idea is to treat each edge of the mesh as an interface of a discontinuous flux problem. Outside the compact of space dependency, our scheme reduces to a standard Godunov method. The contribution of the discontinuous flux theory is that we can build by hand discrete steady states of the scheme, making it, in a way well-balanced.
The next section presents our main result.
2 Description of the scheme and main results
In this section, we describe the construction of our scheme for (1). The novelty lies in the technique used, where after discretizing the space dependency of $H$, at each edge of the mesh, we solve a Riemann problem for a conservation law with discontinuous flux.
Fix a spatial mesh size $\Delta x>0$ and time step $\Delta t>0$. For all $n\in\mathbb{N}$ and $j\in\mathbb{Z}$, set the notations
$t^{n}\coloneq n\Delta t,\quad x_{j}\coloneq j\Delta x,\;x_{j+1/2}\coloneq(j+1/2)\Delta x,\;I_{j}\coloneq\mathopen{]}x_{j-1/2},x_{j+1/2}\mathclose{[}.$
Discretize the initial datum and “the real line”:
$\forall j\in\mathbb{Z},\quad u_{j}^{o}\coloneq\frac{1}{\Delta x}\int_{I_{j}}u_{o}(x)\mathinner{\mathrm{d}{x}},\quad w_{\Delta}(x)\coloneq\sum_{j\in\mathbb{Z}}x_{j}\mathbf{1}_{I_{j}}(x).$
Fix $n\in\mathbb{N}$. Let us explain how, given $(u_{j}^{n})_{j\in\mathbb{Z}}$, we determine $(u_{j}^{n+1})_{j\in\mathbb{Z}}$. Set $\rho^{n}\coloneq\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{I_{j}}$. Let us call $\mathcal{U}^{n}=\mathcal{U}^{n}(t,x)$ the unique entropy solution to following Cauchy problem, set on $\mathopen{]}t^{n},t^{n+1}\mathopen{[}\times\mathbb{R}$:
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl{(}H(w_{\Delta}(x),\mathcal{U}(t,x))\bigr{)}$ $=0$
$\mathcal{U}(t^{n},x)$ $=\rho^{n}(x).$ (2.1)
We recall in Section 3.1 the meaning of solutions to conservation laws with discontinuous flux we use and also establish that (2) admits a unique solution. Notice that since the space dependency is localized in a compact subset of $\mathbb{R}$, outside this compact, (2) reduces to a standard homogeneous conservation law.
Define $u_{j}^{n+1}$ as
$\forall j\in\mathbb{Z},\quad u_{j}^{n+1}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\mathcal{U}^{n}(t^{n+1}-,x)\mathinner{\mathrm{d}{x}}.$ (2.2)
For the numerical analysis, we define the approximate solution $u_{\Delta}$ the following way:
$u_{\Delta}\coloneq\sum_{n=0}^{+\infty}\mathcal{U}^{n}\mathbf{1}_{[t^{n},t^{n+1}[}.$
We can now state our main result.
Theorem 2.1.
Assume that $H$ satisfies (C3)–(CVX) and fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. There exist two constants $\underline{u}$, $\overline{u}$ depending only on $H$ and $u_{o}$ such that under the CFL condition
$2\lambda L\leq 1,\quad\lambda\coloneq\frac{\Delta t}{\Delta x},\;L\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|,$ (2.3)
the following points hold.
  1. (M1)
    The scheme is stable:
[…]
  2. (M2)
    As $(\Delta x,\Delta t)\to(0,0)$, the sequence $(u_{\Delta})_{\Delta}$ converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for all $p\in[1,+\infty\mathclose{[}$ and a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to the unique entropy solution of (1).
The main contribution of the paper is (M1), which proof is detailed in Section 3.2. Indeed, when building a finite volume scheme for a space dependent conservation law, the main difficulty to overcome in the convergence analysis is the obtaining of a priori $\mathbf{L}^{\infty}$ bounds for the approximating sequence. The input of the discontinuous flux theory is that we can build “by hand” discrete steady states of the scheme, making it, in a way, well-balanced.
The proof of (M2) is then somewhat standard, and the details are in Section 3.3. For the convergence analysis, since global $\mathbf{BV}$ bounds are not expected in the context of discontinuous flux, we rely on the compensated compactness method [29, 35], taking advantage of the genuine nonlinearity of $H$ under Assumption (CVX).
3 Proofs
3.1 Discontinuous Flux
In this section, we recall and adapt for our use standard results regarding conservation laws with discontinuous flux. For the purpose of this section, let us fix $f_{l},f_{r}\in\mathbf{C}^{1}(\mathbb{R},\mathbb{R})$ two convex functions satisfying (CVX). Consider:
$\partial_{t}u+\partial_{x}\left(F(x,u)\right)=0,\quad F(x,u)\coloneq\left\{\begin{aligned} f_{l}(u)&\quad\text{if}\;x\leq 0\\ f_{r}(u)&\quad\text{if}\;x>0.\end{aligned}\right.$ (3.1)
Since the works of [1, 2, 3, 4, 13, 19, 24, 26, 37], it is now well known that additionally to the conservation of mass (Rankine-Hugoniot condition)
[…]
an entropy criterion must be imposed at the interface to select one solution. This choice is often guided by physical consideration. Among the most common criteria, we can cite: the minimal jump condition [19, 23, 22], the vanishing viscosity criterion [5], or the flux maximization [3].
If $\Phi_{l}$, resp. $\Phi_{r}$, denotes the Kruzhkov entropy flux associated with $f_{l}$, resp. with $f_{r}$, then in this section, define $\Phi=\Phi(x,u)$ as the Kruzhkov entropy flux associated with $F$:
$\forall x\in\mathbb{R},\;\forall a,b\in\mathbb{R},\quad\Phi(x,a,b)\coloneq\mathop{\rm sgn}(a-b)(F(x,a)-F(x,b))=\left\{\begin{aligned} \Phi_{l}(a,b)&\quad\text{if}\;x\leq 0\\ \Phi_{r}(a,b)&\quad\text{if}\;x>0.\end{aligned}\right.$
For the study of (3.1), we follow [4], where the traces of the solution at the interface $\{x=0\}$ are explicitly treated. It will be useful to have a name for the critical points of $f_{l,r}$, say $\alpha_{l,r}$. Notice that for all $y\in\mathopen{]}\min f_{r},+\infty\mathclose{[}$, the equation $f_{r}(u)=y$ admits exactly two solutions, $S_{r}^{-}(y)<\alpha_{r}<S_{r}^{+}(y)$. When $y=\min f_{r}$, then $S_{r}^{-}(y)=S_{r}^{+}(y)=\alpha_{r}$. This motivates the following definition.
Definition 3.1.
The admissibility germ for (3.1) is the subset $\mathcal{G}$ defined as the union of:
$\mathcal{G}_{1}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\geq\alpha_{l},\;k_{r}=S_{r}^{+}(f_{l}(k_{l}))\right\}:$ (3.2)
$\mathcal{G}_{2}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\leq\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}:$
$\mathcal{G}_{3}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}>\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}.:$
The germ contains all the possible traces along $\{x=0\}$ of the solutions to (3.1). Notice that by construction, any couple in the germ satisfies the Rankine-Hugoniot condition. Conversely, some couples verifying the Rankine-Hugoniot condition have been excluded from the germ, more precisely the ones belonging to
$\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}<\alpha_{l},\;k_{r}=S_{r}^{+}(f_{l}(k_{l}))\right\}.$
The reason lies in the following proposition, see in particular (3.4).
Proposition 3.2.
$\mathcal{G}$ defined in Definition 3.1 is a maximal $\mathbf{L}^{1}$-dissipative germ, meaning that the following properties are satisfied.
(i) For all $(u_{l},u_{r})\in\mathcal{G}$, $f_{l}(u_{l})=f_{r}(u_{r})$.
(ii) Dissipative inequality:
$\forall(u_{l},u_{r}),(k_{l},k_{r})\in\mathcal{G},\quad\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0.$ (3.3)
(iii) Maximality condition: let $(u_{l},u_{r})\in\mathbb{R}^{2}$ such that $f_{l}(u_{l})=f_{r}(u_{r})$. Then
$\forall(k_{l},k_{r})\in\mathcal{G},\quad\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0\implies(u_{l},u_{r})\in\mathcal{G}.$ (3.4)
Proof.  Point (i) is clear, by construction, while (ii) follows from a straightforward case by case study. We now prove (iii) by way of contradiction: let $(u_{l},u_{r})\in\mathbb{R}^{2}$ such that $f_{l}(u_{l})=f_{r}(u_{r})$ and assume that
$\forall(k_{l},k_{r})\in\mathcal{G},\;\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0\quad\text{and}\quad(u_{l},u_{r})\notin\mathcal{G}.$
From the Rankine-Hugoniot condition, we deduce that $u_{r}=S_{r}^{+}(f_{l}(u_{l}))$, with $u_{l}<\alpha_{l}$. Take $k_{r}=S_{r}^{-}(f_{(}k_{l}))$ with $k_{l}\in\mathopen{]}u_{l},\alpha_{l}\mathclose{[}$. Clearly, $(k_{l},k_{r})\in\mathcal{G}_{2}\subset\mathcal{G}$, therefore by assumption, $\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0$. But the choice of $k_{l}$ raises the contradiction:
$\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})=(f_{l}(k_{l})-f_{l}(u_{l}))-(f_{r}(u_{r})-f_{r}(k_{r}))=2(f_{l}(k_{l})-f_{l}(u_{l}))<0.$
$\square$
The theory of [4] proposes an abstract framework for the study of (3.1). On the other hand, the authors of [3, 2] give a formula for the flux at the interface for Riemann problems of (3.1). In the following proposition, we link the two points of view. Let us adopt the notations:
$\forall a,b\in\mathbb{R},\quad a\wedge b\coloneq\min\{a,b\}\quad\text{and}\quad a\vee b\coloneq\max\{a,b\}.$ (3.5)
Proposition 3.3.
For all $(u_{l},u_{r})\in\mathbb{R}^{2}$, define the interface flux:
$F_{\mathrm{int}}(u_{l},u_{r})\coloneq\max\{f_{l}(u_{l}\vee\alpha_{l}),f_{r}(\alpha_{r}\wedge u_{r})\},$ (3.6)
and the remainder term:
$\mathcal{R}(u_{l},u_{r})\coloneq|F_{\mathrm{int}}(u_{l},u_{r})-f_{l}(u_{l})|+|F_{\mathrm{int}}(u_{l},u_{r})-f_{r}(u_{r})|.$ (3.7)
Then the following points hold.
(i) For all $(u_{l},u_{r})\in\mathbb{R}^{2}$, $(u_{l},u_{r})\in\mathcal{G}\iff\mathcal{R}(u_{l},u_{r})=0$.
(ii) For all $(u_{l},u_{r})\in\mathbb{R}^{2}$ and for all $(k_{l},k_{r})\in\mathcal{G}$, $\Phi_{r}(u_{r},k_{r})-\Phi_{l}(u_{l},k_{l})\leq\mathcal{R}(u_{l},u_{r})$.
Proof.  The fact that we have an “explicit” description of $\mathcal{G}$ facilitates the proof.
(i) If $(u_{l},u_{r})\in\mathcal{G}_{1}\cup\mathcal{G}_{2}$, then $u_{l}-\alpha_{l}$ and $u_{r}-\alpha_{r}$ have the same sign, therefore
$F_{\mathrm{int}}(u_{l},u_{r})=\left\{\begin{aligned} f_{l}(u_{l})&\quad\text{if}\;u_{l}\geq\alpha_{l}\\ f_{r}(u_{r})&\quad\text{if}\;u_{l}<\alpha_{l}\end{aligned}\right.=f_{l}(u_{l})=f_{r}(u_{r}).$
On the other hand, if $(u_{l},u_{r})\in\mathcal{G}_{3}$, then
$F_{\mathrm{int}}(u_{l},u_{r})=\max\{f_{l}(u_{l}),f_{r}(u_{r})\}=f_{l}(u_{l})=f_{r}(u_{r}).$
This ensures that $\mathcal{R}(u_{l},u_{r})=0$.
Conversely, assume that $\mathcal{R}(u_{l},u_{r})=0$. In particular, the Rankine-Hugoniot holds. We deduce that
$\mathcal{R}(u_{l},u_{r})=0\implies(u_{l},u_{r})\in\mathcal{G}\;\;\text{or}\;\;u_{r}=S_{r}^{+}(f_{l}(u_{l})),\;u_{l}<\alpha_{l}.$
But in the latter case, $F_{\mathrm{int}}(u_{l},u_{r})=\max\{\min f_{l},\min f_{r}\}\neq f_{l}(u_{l})$. Therefore, $(u_{l},u_{r})\in\mathcal{G}$.
(ii) For clarity, set $\mu\coloneq\max\{\min f_{l},\min f_{r}\}$ and notice that
$\mathcal{R}(u_{l},u_{r})=\left\{\begin{aligned} |f_{l}(u_{l})-\mu|+|f_{r}(u_{r})-\mu|&\quad\text{if}\;u_{l}<\alpha_{l}\;\text{and}\;u_{r}\geq\alpha_{r}\\ |f_{l}(u_{l})-f_{r}(u_{r})|&\quad\text{otherwise.}\end{aligned}\right.$
Let us call $q\coloneq\Phi_{r}(u_{r},k_{r})-\Phi_{l}(u_{l},k_{l})$. We proceed with a case by case study. We use the fact that $f_{l,r}$ is decreasing on $\mathopen{]}-\infty,\alpha_{l,r}]$ and increasing on $[\alpha_{l,r},+\infty\mathclose{[}$.
Case 1: $u_{l}\leq k_{l}$ and $u_{r}\leq k_{r}$. Then $q=f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r})$.
Case 2: $u_{l}>k_{l}$ and $u_{r}>k_{r}$. Similar to Case 1.
Case 3: $u_{l}>k_{l}$ and $u_{r}\leq k_{r}$. Then $q=(f_{r}(k_{r})-f_{r}(u_{r}))+(f_{l}(k_{l})-f_{l}(u_{l}))$. Notice that because of the germ structure, the case $k_{l}<u_{l}<\alpha_{l}$ and $\alpha_{r}\leq u_{r}\leq k_{r}$ cannot happen. Therefore, in Case 3, $\mathcal{R}(u_{l},u_{r})=|f_{l}(u_{l})-f_{r}(u_{r})|$. We have
$\bullet$ $k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}$
$\implies q\leq f_{l}(k_{l})-f_{l}(u_{l})=f_{r}(u_{r})-f_{l}(u_{l})+(f_{r}(k_{r})-f_{r}(u_{r}))\leq f_{r}(u_{r})-f_{l}(u_{l})\leq\mathcal{R}(u_{l},u_{r})$
$k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}<k_{r}\;\text{does not happen}$
$\bullet$ $k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}\;\text{does not happen}$
$\bullet$ $\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}\implies q\leq 0$
$k_{l}<\alpha_{l}<u_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}$
$\implies q\leq f_{l}(k_{l})-f_{l}(u_{l})=f_{r}(u_{r})-f_{l}(u_{l})+(f_{r}(k_{r})-f_{r}(u_{r}))\leq f_{r}(u_{r})-f_{l}(u_{l})\leq\mathcal{R}(u_{l},u_{r})$
$\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\leq k_{r}$
$\implies q\leq f_{r}(k_{r})-f_{r}(u_{r})=f_{l}(u_{l})-f_{r}(u_{r})+(f_{l}(k_{l})-f_{l}(u_{l}))\leq f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r})$
$\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\leq k_{r}\;\text{does not happen}$
$\bullet$ $\alpha_{l}<k_{l}<u_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}$
$\implies q\leq f_{r}(k_{r})-f_{r}(u_{r})=f_{l}(u_{l})-f_{r}(u_{r})+(f_{l}(k_{l})-f_{l}(u_{l}))\leq f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r})$
$k_{l}<\alpha_{l}<u_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}\;\text{does not happen.}$.
When we say “does not happen”, we mean that germ geometry prevents the situation from occurring.
Case 4: $u_{l}\leq k_{l}$ and $u_{r}>k_{r}$. Similar to Case 3.   $\square$
Remark 3.1.
Notice that for all $k\in\mathbb{R}$,
$\mathcal{R}(k,k)=\left\{\begin{aligned} |\max\{\min f_{l},\min f_{r}\}-f_{l}(k)|+|\max\{\min f_{l},\min f_{r}\}-f_{r}(k)|&\quad\text{if}\;\alpha_{r}\leq k<\alpha_{l}\\ |f_{l}(k)-f_{r}(k)|&\quad\text{otherwise.}\end{aligned}\right.$
We are now in position to properly give the definition of solutions for (3.1).
Proposition 3.4.
Define $\mathcal{G}$ as (3.2) and $\mathcal{R}$ as (3.7). Let $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. Then the following statements are equivalent.
1. For all test functions $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and for all $(k_{l},k_{r})\in\mathbb{R}^{2}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-\kappa(x)|\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}|u_{o}(x)-\kappa(x)|\varphi(0,x)\mathinner{\mathrm{d}{x}}$
$+\int_{0}^{+\infty}\mathcal{R}(k_{l},k_{r})\varphi(t,0)\mathinner{\mathrm{d}{t}}\geq 0.$ (3.8)
2.(i) For all test functions $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R}^{*},\mathbb{R}^{+})$ and for all $(k_{l},k_{r})\in\mathbb{R}^{2}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-\kappa(x)|\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}|u_{o}(x)-\kappa(x)|\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (3.9)
2.(ii) For a.e. $t\in\mathopen{]}0,+\infty\mathclose{[}$, $(u(t,0-),u(t,0+))\in\mathcal{G}$.
When one of these statements holds, we say that $u$ is an entropy solution to (3.1) with initial datum $u_{o}$.
Proof.  See the proof of [4, Theorem 3.18].   $\square$
Remark 3.2.
In both (3.4) and (3.9), $\kappa$ denotes the piecewise constant function
$\kappa(x)=\left\{\begin{aligned} k_{l}&\quad\text{if}\;x\leq 0\\ k_{r}&\quad\text{if}\;x>0.\end{aligned}\right.$
In the second definition, the traces of the solution along $\{x=0\}$ explicitly appear. The condition
[…]
is to be understood in the sense of strong traces. An important fact we stress is that it is not restrictive to assume that general entropy solutions, that is to say bounded functions verifying (3.9), admit strong traces. Usually, it is ensured provided a nondegeneracy assumption on the flux function. Here, since $f_{l}$ and $f_{r}$ are genuinely nonlinear in the sense that
$\forall s\in\mathbb{R},\quad\text{meas}\biggl{(}\left\{p\in\mathbb{R}\;:\;f_{l,r}^{\prime}(p)=s\right\}\biggr{)}=0,$
existence of strong traces is well known since [38], see also [33].
We now state the well-posedness result regarding (3.1).
Theorem 3.5.
Let $f_{l},f_{r}\in\mathbf{C}^{1}(\mathbb{R},\mathbb{R})$ verify (CVX).
  1. (D1)
    Fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. Then there exist a unique entropy solution to (3.1) with initial datum $u_{o}$.
  2. (D2)
    Let $u_{o},v_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We denote by $u$, resp. $v$, an entropy solution to (3.1) with initial datum $u_{o}$, resp. $v_{o}$. Then, there exists $L>0$ such that for all $R>0$ and for all $t>0$,
    $\int_{|x|\leq R}|u(t,x)-v(t,x)|\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}|u_{o}(x)-v_{o}(x)|\mathinner{\mathrm{d}{x}}$ (3.10)
    $\int_{|x|\leq R}(u(t,x)-v(t,x))^{+}\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}(u_{o}(x)-v_{o}(x))^{+}\mathinner{\mathrm{d}{x}}.$
Proof.  The proofs of the stability bounds can be found in [4, Theorems 3.11-3.19]. Regarding the existence, one can build a simple finite volume scheme and prove its convergence, see for instance [2]. The flux to put at the interface $\{x=0\}$ in the scheme is precisely $F_{\mathrm{int}}$.   $\square$
We conclude this section explaining how to solve the Riemann problem for (3.1). Indeed, the structure of the solution to Riemann problems will be used in the convergence analysis, see in particular Section 3.3.1.
Given $(u_{l},u_{r})\in\mathbb{R}^{2}$, we solve (3.1) with the initial condition
$u_{o}(x)=\left\{\begin{array}[]{ccl}u_{l}&\text{if}&x<0\\ u_{r}&\text{if}&x>0.\end{array}\right.$
First, compute the interface flux given by (3.6). We have:
$F_{\mathrm{int}}(u_{l},u_{r})=\left\{\begin{array}[]{ccl}\max\{f_{l}(\alpha_{l}),f_{r}(u_{r})\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(I)}\\ \max\{\min f_{l},\min f_{r}\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(II)}\\ \max\{f_{l}(u_{l}),f_{r}(u_{r})\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(III)}\\ \max\{f_{l}(u_{l}),\min f_{r}\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(IV)}\end{array}\right.$ (3.11)
The value of $F_{\mathrm{int}}(u_{l},u_{r})$ imposes the value of one of the traces at the interface. For instance, if $F_{\mathrm{int}}(u_{l},u_{r})=f_{r}(u_{r})$, then the right trace of the solution along $\{x=0\}$ is equal to $u_{r}$. To determine the left trace, one solves $f_{l}(\gamma)=f_{r}(u_{r})$, $\gamma\in\mathbb{R}$, with the requirement that the solution to the classical Riemann problem with flux $f_{l}$ and states $(\gamma,u_{r})$ only displays waves of negative speeds. We list below the different structure of the solution.
If $(u_{l},u_{r})\in\mathcal{G}$,
then the solution is a stationary non-classical shock (SNS).
Case (I)
The solution is the gluing of shock/rarefaction - SNS - constant $u_{r}$ or rarefaction - SNS - shock.
Case (II)
The solution is the gluing of shock/rarefaction - SNS - rarefaction, or rarefaction - SNS - shock/rarefaction.
Case (III)
The solution is the gluing of shock - SNS - constant $u_{r}$, or constant $u_{l}$ SNS - shock.
Case (IV)
The solution is the gluing of shock - SNS - rarefaction, or constant $u_{l}$ SNS - shock/rarefaction.
In any case, the solution presents at most one stationary non-classical shock and at most one classical shock.
The content of this section can be adapted in a straightforward way to the case where $F$ in (3.1) presents a finite number of space discontinuities.
3.2 Steady states and stability
This section is devoted to the proof of the $\mathbf{L}^{\infty}$ stability of the scheme.
Lemma 3.6.
Assume that $H$ satisfies (C3)–(CVX). Then, there exists a unique function $\alpha\in\mathbf{C}^{2}(\mathbb{R},\mathbb{R})$ such that for all $x\in\mathbb{R}$, $\partial_{u}H(x,\alpha(x))=0$. Moreover,
$\forall x\in\mathbb{R},\quad|x|\geq X\implies\alpha^{\prime}(x)=0.$
Proof.  Straightforward application of the Implicit Function Theorem.   $\square$
Introduce $J\in\mathbb{N}^{*}$ such that $X\in I_{J}$. Recall that the space dependency of $H$ is localized in $[-X,X]$.
Set for all $j\in\mathbb{Z}$, $\alpha_{j}\coloneq\alpha(x_{j})$ and notice that
$j\geq J+1\implies\alpha_{j}=\alpha(X)\quad\text{and}\quad j\leq-J-1\implies\alpha_{j}=\alpha(-X).$
It will be convenient to adopt the notation
$\forall j\in\mathbb{Z},\quad h_{j}(u)\coloneq H(x_{j},u).$
Let us precise that for sufficiently small time interval $\Delta t$, the waves emanating at each interface $x=x_{j+1/2}$ do not interact. Consequently, (2.2) rewrites as
$\forall j\in\mathbb{Z},\quad u_{j}^{n+1}=u_{j}^{n}-\lambda(F_{\mathrm{int}}^{j+1/2}(u_{j}^{n},u_{j+1}^{n})-F_{\mathrm{int}}^{j-1/2}(u_{j-1}^{n},u_{j}^{n})),$ (3.12)
where following Section 3.1,
$F_{\mathrm{int}}^{j+1/2}(a,b)\coloneq\max\biggl{\{}\mathrm{god}_{j}(a,\alpha_{j}),\;\mathrm{god}_{j+1}(\alpha_{j+1},b)\biggr{\}}$
is the flux across the interface $x=x_{j+1/2}$, with $\mathrm{god}_{j}$ denoting the Godunov flux associated with $h_{j}$.
It is worth mentioning that for all $j\in\mathbb{Z}$, $F_{\mathrm{int}}^{j+1/2}$ is a monotone, locally Lipschitz, numerical flux, see [2, Section 4]. Also, let us precise that in general, $F_{\mathrm{int}}^{j+1/2}$ is not consistent, meaning that for all $k\in\mathbb{R}$, $F_{\mathrm{int}}^{j+1/2}(k,k)$ is not equal to $h_{j}(k)$ or to $h_{j+1}(k)$. However, using the regularity of $H$, we can prove that a consistency holds, at the limit.
Lemma 3.7.
Assume that $H$ satisfies (C3)–(CVX). Then for any $k\in\mathbb{R}$, there exists a positive constant $\mu_{k}$ such that for all $j\in\mathbb{Z}$,
$\left|F_{\mathrm{int}}^{j+1/2}(k,k)-h_{j+1}(k)\right|\leq\mu_{k}\;\Delta x.$
Proof.  Let $j\in\mathbb{Z}$. Recall that for all $k\in\mathbb{R}$,
$F_{\mathrm{int}}^{j+1/2}(k,k)=\left\{\begin{array}[]{cclc}\max\{h_{j}(\alpha_{j}),h_{j+1}(k)\}&\text{if}&k\leq\alpha_{j},\alpha_{j+1}&\text{(I)}\\ \max\{h_{j}(\alpha_{j}),h_{j+1}(\alpha_{j+1})\}&\text{if}&\alpha_{j+1}\leq k\leq\alpha_{j}&\text{(II)}\\ \max\{h_{j}(k),h_{j+1}(k)\}&\text{if}&\alpha_{j}\leq k\leq\alpha_{j+1}&\text{(III)}\\ \max\{h_{j}(k),h_{j+1}(\alpha_{j+1})\}&\text{if}&k>\alpha_{j},\alpha_{j+1}&\text{(IV)}\end{array}\right.$
The proof reduces to a case by case study. Set $\delta\coloneq\left|F_{\mathrm{int}}^{j+1/2}(k,k)-h_{j+1}(k)\right|$
Case (I) If $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j+1}(k)$, then $\delta=0$. Otherwise, $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j}(\alpha_{j})$, meaning that $h_{j}(\alpha_{j})\geq h_{j+1}(k)$. Therefore,
$\delta=h_{j}(\alpha_{j})-h_{j+1}(k)\leq h_{j}(k)-h_{j+1}(k)\leq\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x.$
Case (II) In that case, the estimate follows from the fact that
$|h_{j}(\alpha_{j})-h_{j+1}(k)|\leq\left(\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{|p|\leq\|\alpha\|_{\mathbf{L}^{\infty}}}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\Delta x.$
Case (III) In that case, we can bound $\delta$ by $\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x$.
Case (IV) If $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j}(k)$, then we bound $\delta$ by $\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x$. Otherwise, $F_{\mathrm{int}}^{j+1/2}(k,k)=h_{j+1}(\alpha_{j+1})$, meaning that $h_{j+1}(\alpha_{j+1})\geq h_{j}(k)$. Therefore,
$\delta$ $\leq|h_{j+1}(\alpha_{j+1})-h_{j}(k)|+|h_{j}(k)-h_{j+1}(k)|$
$=(h_{j+1}(\alpha_{j+1})-h_{j}(k))+|h_{j}(k)-h_{j+1}(k)|$
$\leq 2|h_{j}(k)-h_{j+1}(k)|$
$\leq 2\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\;\Delta x.$
$\square$
To prove (M1), we explicitly construct steady states of the scheme (3.12).
Definition 3.8.
We say that a sequence $(v_{j})_{j}$ is a steady state of (3.12) if
$\forall j\in\mathbb{Z},\quad F_{\mathrm{int}}^{j-1/2}(v_{j-1},v_{j})=F_{\mathrm{int}}^{j+1/2}(v_{j},v_{j+1}).$
Following Section 3.1, it will be useful to introduce for all $j\in\mathbb{Z}$ and $y\in\mathopen{]}\min h_{j},+\infty\mathclose{[}$, the two solutions $S_{j}^{-}(y)<\alpha_{j}<S_{j}^{+}(y)$ of the equation $h_{j}(x)=y$, $x\in\mathbb{R}$.
Lemma 3.9.
Assume that $H$ satisfies (C3)–(CVX). Let $c\in\mathbb{R}$ such that $c\geq\sup_{\mathbb{R}}\alpha$. Define for all $j\in\mathbb{Z}$,
$v_{j+1}\coloneq\left\{\begin{array}[]{ccl}c&\text{if}&j\leq-J-2\\ S_{j+1}^{+}(h_{j}(v_{j}))&\text{if}&j\geq-J-1,\end{array}\right.\quad\quad w_{j}\coloneq\left\{\begin{array}[]{ccl}S_{j}^{+}(h_{j+1}(w_{j+1}))&\text{if}&j\leq J\\ c&\text{if}&j\geq J+1.\end{array}\right.$
Then $(v_{j})_{j}$ and $(w_{j})_{j}$ are steady states of (3.12) in the sense of Definition 3.8.
Proof.  The key point is that for all $j\in\mathbb{Z}$, $(v_{j},v_{j+1})$ and $(w_{j},w_{j+1})$ are chosen in the germ $\mathcal{G}_{j+1/2}$, see Definition 3.1, which ensures by Proposition 3.3 that
$h_{j}(v_{j})=F_{\mathrm{int}}^{j+1/2}(v_{j},v_{j+1})=h_{j+1}(v_{j+1}).$
$\square$
Remark 3.3.
Keep the notations of Lemma 3.9.
(i) Notice that $(v_{j})_{j}$ and $(w_{j})_{j}$ are stationary. For instance, for all integers $j\geq J+2$,
$h_{j-1}(v_{j-1})=h_{j}(v_{j})=h_{j-1}(v_{j})\implies v_{j-1}=v_{j}.$
(ii) For all $c\in\mathbb{R}$ such that $c\leq\inf_{\mathbb{R}}\alpha$, the two following sequences are steady states as well:
$v_{j+1}=\left\{\begin{array}[]{ccl}c&\text{if}&j\leq-J-2\\ S_{j+1}^{-}(h_{j}(v_{j}))&\text{if}&j\geq-J-1,\end{array}\right.\quad\quad w_{j}=\left\{\begin{array}[]{ccl}S_{j}^{-}(h_{j+1}(w_{j}))&\text{if}&j\leq J\\ c&\text{if}&j\geq J+1.\end{array}\right.$
Lemma 3.10.
Assume that $H$ satisfies (C3)–(CVX). Let $(m,M)\in\mathbb{R}^{2}$ such that $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$. Then there exist two steady states of (3.12), $(m_{j})_{j}$ and $(M_{j})_{j}$, such that:
(i) $(m_{j})_{j}$ is bounded by below ;
(ii) $(M_{j})_{j}$ is bounded by above ;
(iii) for all $j\in\mathbb{Z}$, $m_{j}\leq m$ and $M_{j}\geq M$.
Proof.  We only give the details for the construction of $(M_{j})_{j}$.
Set $\mathcal{U}_{h}\coloneq\max_{x\in\mathbb{R}}H(x,M)$. By monotonicity, for all $j\in\mathbb{Z}$, $S_{j}^{+}(\mathcal{U}_{h})\geq M$. Therefore, if $\overline{M}\coloneq\max_{j\in\mathbb{Z}}S_{j}^{+}(\mathcal{U}_{h})$, then $\overline{M}\geq M\geq\sup_{x\in\mathbb{R}}\alpha(x)$. Define
$\forall j\in\mathbb{Z},\quad M_{j+1}\coloneq\left\{\begin{array}[]{ccl}\overline{M}&\text{if}&j\leq-J-2\\ S_{j+1}^{+}(h_{j}(M_{j}))&\text{if}&j\geq-J-1.\end{array}\right.$
Clearly, $(M_{j})_{j}$ is a steady state of the scheme.
Claim 1:
for all $j\in\mathbb{Z}$, $M_{j}\geq M$. Indeed, for all $j\in\mathbb{Z}$,
$h_{j}(M_{j})=h_{j+1}(M_{j+1})=h_{-J-1}(\max_{i\in\mathbb{Z}}S_{i}^{+}(\mathcal{U}_{h}))\geq h_{-J-1}(S_{-J-1}^{+}(\mathcal{U}_{h}))=\mathcal{U}_{h}\geq h_{j}(M),$
from which we deduce, by monotonicity, that $M_{j}\geq M$. The claim is proved.
Claim 2:
$(M_{j})_{j}$ is bounded by above. By definition, for all $j\in\mathbb{Z}$,
$M_{j}=S_{j}^{+}\left(H(-X,\overline{M})\right)\leq\max_{x\in\mathbb{R}}h_{+}^{-1}\left(x,H(-X,\overline{M})\right)\coloneq\overline{u},$
where for all $x\in\mathbb{R}$, we denote by $u\mapsto h_{+}^{-1}(x,u)$ the reciprocal of $u\mapsto h(x,u)$ on $[\alpha(x),+\infty\mathclose{[}$. We see that $\overline{u}$ only depends on $H$ and $M$.   $\square$
Proof of (M1).  With reference to Lemma 3.6, fix $(m,M)\in\mathbb{R}^{2}$ satisfying $m\leq\inf_{\mathbb{R}}\alpha$ and $M\geq\sup_{\mathbb{R}}\alpha$, such that for a.e. $x\in\mathbb{R}$, $m\leq u_{o}(x)\leq M$. Consider $(m_{j})_{j}$ and $(M_{j})_{j}$, steady states given by Lemma 3.10.
Define the piecewise constant functions
$(m_{\Delta},M_{\Delta})\coloneq\sum_{j\in\mathbb{Z}}(m_{j},M_{j})\mathbf{1}_{I_{j}},$
and $\underline{u}$, $\overline{u}$ as they were in the proof of Lemma 3.10.
The key point is that $m_{\Delta}$ and $M_{\Delta}$ are both stationary solutions to
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl{(}H(w_{\Delta}(x),\mathcal{U}(t,x))\bigr{)}=0.$
Therefore, for all $n\in\mathbb{N}$, (D2) of Theorem 3.5 ensures that
$m_{\Delta}\leq\rho^{n}\leq M_{\Delta}\implies\forall t\in[t^{n},t^{n+1}],\quad m_{\Delta}\leq\mathcal{U}^{n}(t)\leq M_{\Delta}.$
Since by construction of the sequences $(m_{j})_{j\in\mathbb{Z}}$ and $(M_{j})_{j\in\mathbb{Z}}$, we have $m_{\Delta}\leq\rho^{0}\leq M_{\Delta}$, a direct induction provides:
$\forall t\in\mathbb{R}^{+},\quad m_{\Delta}\leq u_{\Delta}(t)\leq M_{\Delta},$
immediately leading to (2.4) by definition of $\underline{u}$ and $\overline{u}$.
The proof is concluded.   $\square$
3.3 Convergence of the scheme
We first establish the strong compactness of $(u_{\Delta})_{\Delta}$ with a compensated compactness argument.
3.3.1 Compensated compactness
The compensated compactness method and its applications to systems of conservation laws is for instance reviewed in [10, 29]. Modified for our use, the compensated compactness lemma reads as follows.
Lemma 3.11.
Assume that $H$ satisfies (CH)–(CVX), ensuring its genuine nonlinearity in the sense of
$\forall(x,s)\in\mathbb{R}^{2},\quad\text{meas}\biggl{(}\left\{u\in\mathbb{R}\;:\;\partial_{u}H(x,u)=s\right\}\biggr{)}=0.$ (3.13)
Let $(u_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ be a bounded sequence of $\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ such that for all $k\in\mathbb{R}$ and for any $i\in\{1,2\}$, the sequence $(\partial_{t}S_{i}(u_{\mathrm{\varepsilon}})+\partial_{x}Q_{i}(x,u_{\mathrm{\varepsilon}}))_{\mathrm{\varepsilon}}$ belongs to a compact subset of $\mathbf{H}_{\mathbf{loc}}^{-1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$, where
$\begin{array}[]{cc}S_{1}(u)\coloneq u-k&Q_{1}(x,u)\coloneq H(x,u)-H(x,k)\\[5.0pt] S_{2}(u)\coloneq H(x,u)-H(x,k)&\displaystyle{Q_{2}(x,u)\coloneq\int_{k}^{u}\partial_{u}H(x,\xi)^{2}\mathinner{\mathrm{d}{\xi}}}.\end{array}$ (3.14)
Then there exists a subsequence of $(u_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ that converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for all $p\in[1,+\infty\mathclose{[}$ and a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to some function $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
To prove the $\mathbf{H}_{\mathbf{loc}}^{-1}$ compactness required in Lemma 3.11 we will use the following technical result, see [32, 15, 35].
Lemma 3.12.
Let $q,r\in\mathbb{R}$ such that $1<q<2<r$. Let $(\mu_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ be a sequence of distributions such that:
(i) $(\mu_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ belongs to a compact subset of $\mathbf{W}_{\mathbf{loc}}^{-1,q}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
(ii) $(\mu_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ is bounded in $\mathbf{W}_{\mathbf{loc}}^{-1,r}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
Then $(\mu_{\mathrm{\varepsilon}})_{\mathrm{\varepsilon}}$ belongs to a compact subset of $\mathbf{H}_{\mathbf{loc}}^{-1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
For the compactness analysis, we take inspiration from [22, 23].
Let $(S,Q)$ be an entropy/entropy flux pair associated with $H$. The entropy dissipation of $u_{\Delta}$ is defined as
$E_{\Delta}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S(u_{\Delta})\partial_{t}\phi+Q(w_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}).$ (3.15)
Fix $n\in\mathbb{N}$ and $j\in\mathbb{Z}$. Let us consider the entropy dissipation in $P_{j+1/2}^{n}\coloneq[t^{n},t^{n+1}[\times]x_{j},x_{j+1}[$. By integration by parts and the fact that $u_{\Delta}$ is the exact solution of a Riemann problem in $P_{j+1/2}^{n}$, we write:
$\iint_{P_{j+1/2}^{n}}$ $S(u_{\Delta})\partial_{t}\phi+Q(w_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$=\int_{x_{j}}^{x_{j+1}}S(u_{\Delta}(t^{n+1}-,x))\phi(t^{n+1},x)-S(u_{\Delta}(t^{n},x))\phi(t^{n},x)\mathinner{\mathrm{d}{x}}$
$+\int_{t^{n}}^{t^{n+1}}Q(x_{j+1},u_{j+1}^{n})\phi(t,x_{j+1})-Q(x_{j},u_{j}^{n})\phi(t,x_{j})\mathinner{\mathrm{d}{t}}$
$+\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}$
$+\int_{t^{n}}^{t^{n+1}}\biggl{(}Q(x_{j},u_{\Delta}(t,x_{j+1/2}-))-Q(x_{j+1},u_{\Delta}(t,x_{j+1/2}+))\biggr{)}\phi(t,x_{j+1/2})\mathinner{\mathrm{d}{t}}.$
With reference to Section 3.1, in the right-hand side of the previous equality:
$\bullet$ The third integral is the contribution of the (eventual) classical shock in the solution. Hence, $\sigma$ is the shock velocity, given by the Rankine-Hugoniot condition, $y$ is the shock curve and
$[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\coloneq(\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta}))(t,y(t)+)-(\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta}))(t,y(t)-).$
It is relevant to say that $w_{\Delta}$ is continuous along $\{x=y(t)\}$, equal to either $x_{j}$ or $x_{j+1}$, depending on the sign of $\sigma$.
$\bullet$ The last integral is the contribution of the (eventual) stationary non-classical shock in the solution.
Taking the sum for $n\in\mathbb{N}$ and $j\in\mathbb{Z}$, we see that the entropy dissipation rewrites as
$\begin{array}[]{clr}E_{\Delta}(\phi)&\displaystyle{=-\int_{\mathbb{R}}S(u_{\Delta}(0,x))\phi(0,x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{1}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=1}^{+\infty}\int_{\mathbb{R}}(S(u_{\Delta}(t^{n}-,x))-S(u_{\Delta}(t^{n},x)))\phi(t^{n},x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{2}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{3}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{t^{n}}^{t^{n+1}}\biggl{(}Q(x_{j},u_{\Delta}(t,x_{j+1/2}-))-Q(x_{j+1},u_{\Delta}(t,x_{j+1/2}+))\biggr{)}\phi(t,x_{j+1/2})\;\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{4}(\phi).\end{array}$ (3.16)
As we previously explained, for all $n\in\mathbb{N}$ and $j\in\mathbb{Z}$, the sum $\sum_{y}$ in (3.16) is either over an empty set, or over a singleton.
First, let us give a bound on the variation of the approximate solution across the discrete time levels.
Lemma 3.13.
Assume that $H$ satisfies (CH)–(CVX). Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists a constant $c_{1}>0$ depending on $T$, $R$, $u_{o}$, and $H$ such that
$\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|u_{\Delta}(t^{n},x)-u_{\Delta}(t^{n}-,x)|^{2}\mathinner{\mathrm{d}{x}}\leq c_{1}$ (3.17)
and $\sum_{n=0}^{N}\sum_{|j|\leq K}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\mathinner{\mathrm{d}{t}}\leq c_{1}.$
Proof.  For the purpose of this proof, let us choose $S(u)=\frac{u^{2}}{2}$ and $Q=Q(x,u)$ its entropy flux. Let $(\phi_{k})_{k}$ be a sequence of nonnegative test functions that converges to $\phi\coloneq\mathbf{1}_{[0,T]\times[-R,R]}$. At the limit $k\to+\infty$ in (3.16), we obtain:
$I_{2}(\phi)+I_{3}(\phi)=\int_{|x|\leq R}|u_{\Delta}(0,x)|^{2}\mathinner{\mathrm{d}{x}}-I_{4}(\phi)\leq 2R\|u_{o}\|_{\mathbf{L}^{\infty}}^{2}-I_{4}(\phi).$
Note that $I_{3}(\phi)$ is nonnegative because the (eventual) shock is classical and therefore produces entropy.
Then, notice that by definition of $(u_{j}^{n})_{j,n}$,
$I_{2}(\phi)$ $=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}u_{\Delta}(t^{n}-,x)^{2}-u_{\Delta}(t^{n},x)^{2}\;\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\left\{\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}+2u_{\Delta}(t^{n},x)(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))\right\}\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}\mathinner{\mathrm{d}{x}}+\sum_{n=1}^{N}\sum_{|j|\leq K}u_{j}\underbrace{\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})\mathinner{\mathrm{d}{x}}}_{=0}.$
Finally, as a consequence of Proposition 3.3 (ii), for all $k\in[\underline{u},\overline{u}]$,
$\Phi(x_{j+1/2}-,u_{\Delta}(t,x_{j+1/2}-),k)-\Phi(x_{j+1/2}+,u_{\Delta}(t,x_{j+1/2}+),k)\geq-\mathcal{R}_{j+1/2}(k,k).$
Using Remark 3.1, we can bound this remainder term as
$\mathcal{R}_{j+1/2}(k,k)\leq\left(2\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\right)\Delta x.$
Using an approximation argument to pass from Kruzhkov entropies to any entropy, see [22], we obtain
$I_{4}(\phi)\geq-2TR\left(2\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\right).$
Estimate (3.17) follows by putting the bounds on $I_{1}(\phi)$ and $I_{4}(\phi)$ together.   $\square$
We can convert (3.17) into an estimate on the spatial variation of the approximate solutions, following the arguments of [16, Page 67]. For the sake of clarity, we set $\rho_{j+1/2}^{n,\pm}\coloneq u_{\Delta}(t^{n},x_{j+1/2}\pm)$.
Lemma 3.14.
Assume that $H$ satisfies (CH)–(CVX). Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists a constant $c_{2}>0$ depending on $T$, $R$, $u_{o}$ and $H$ such that
$\sum_{n=0}^{N}\sum_{|j|\leq K}\left\{(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+(u_{j}^{n}-\rho_{j+1/2}^{n,-})^{2}+(u_{j+1}^{n}-\rho_{j+1/2}^{n,+})^{2}\right\}\Delta x\leq c_{2}.$ (3.18)
We are now in position to prove the $\mathbf{H}_{\mathbf{loc}}^{-1}$ compactness of the sequence of measures that defines the entropy dissipation.
Lemma 3.15.
Assume that $H$ satisfies (CH)–(CVX). Let $(S_{i},Q_{i})_{i\in\{1,2\}}$ be the entropy/entropy flux pairs defined in Lemma 3.11. Then for any $i\in\{1,2\}$, the sequence of distributions $\mu_{i}$ defined by
$\mu_{i}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S_{i}(u_{\Delta})\partial_{t}\phi+Q_{i}(x,u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R})$
belongs to a compact subset of $\mathbf{H}_{\mathbf{loc}}^{-1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
Proof.  We follow the proofs of [22, Lemma 5.5] or [23, Lemma 4.5]. To start, let us fix $(S,Q)$ a smooth entropy/entropy flux pair.
Let us fix a bounded open subset $\Omega\subset\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$, say $\Omega\subset[0,T]\times[-R,R]$ for some $T>0$ and $R>0$. Call $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}[$ and $R\in I_{K}$, and finally, let $\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\Omega,\mathbb{R})$. We split $\mu$ as
$\mu(\phi)=\int_{0}^{+\infty}\int_{\mathbb{R}}(Q(x,u_{\Delta})-Q(w_{\Delta},u_{\Delta}))\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+E_{\Delta}(\phi),$
where $E_{\Delta}(\phi)$ is given by (3.15).
The definition of $w_{\Delta}$ and regularity of $H$ ensures that for any $q_{1}\in\mathopen{]}1,2]$, the first term of the right-hand side is strongly compact in $\mathbf{W}^{-1,q_{1}}(\Omega,\mathbb{R})$. We now estimate $E_{\Delta}(\phi)$.
First, observe that since $(u_{\Delta})_{\Delta}$ is bounded in $\mathbf{L}^{\infty}(\Omega,\mathbb{R})$, we have
$|E_{\Delta}(\phi)|\leq\biggl{(}\sup_{\underline{u}\leq p\leq\overline{u}}|S(p)|+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|Q(x,p)|\biggr{)}\cdot\|\phi\|_{\mathbf{W}^{1,\infty}},$
implying that $(E_{\Delta}(\phi))_{\Delta}$ is bounded in $\mathbf{W}_{\mathbf{loc}}^{-1,r}(\Omega,\mathbb{R})$ for any $r\in\mathopen{]}2,+\infty\mathclose{[}$.
Keeping the notations of (3.16), we bound $I_{1}(\phi)$ and $I_{3}(\phi)$ as
$|I_{1}(\phi)|\leq\|S(u_{o})\|_{\mathbf{L}^{1}([-R,R])}\;\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},\quad I_{3}(\phi)\leq c_{1}\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
where we used Lemma 3.13 for $I_{3}(\phi)$. Consequently, $I_{1}(\phi)$ and $I_{3}(\phi)$ are bounded in the space $\mathcal{M}(\Omega,\mathbb{R})$ of bounded Radon measures, which is compactly embedded in $\mathbf{W}^{-1,q_{2}}(\Omega,\mathbb{R})$ if $q_{2}\in\mathopen{]}1,2\mathclose{[}$. Therefore, $I_{1}(\phi)$ and $I_{3}(\phi)$ belong to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega,\mathbb{R})$.
Now, to estimate $I_{2}(\phi)$, split it as $I_{2}(\phi)=I_{2,1}(\phi)+I_{2,2}(\phi)$ where
$I_{2,1}(\phi)$ :-$\coloneq\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))\phi(t^{n},x_{j})\mathinner{\mathrm{d}{x}}:$
$I_{2,2}(\phi)$ :-$\coloneq\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))(\phi(t^{n},x)-\phi(t^{n},x_{j}))\mathinner{\mathrm{d}{x}}.:$
We handle $I_{2,1}(\phi)$ by introducing an intermediary point $\omega_{j}^{n}$ such that
$S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})=S^{\prime}(u_{j}^{n})(u_{\Delta}(t^{n}-,x)-u_{j}^{n})+\frac{S^{\prime\prime}(\omega_{j}^{n})}{2}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})^{2}.$
Taking into account the definition of $u_{j}^{n}$ and Lemma 3.14, we obtain
$|I_{2,1}(\phi)|\leq\frac{c_{1}}{2}\sup_{\underline{u}\leq p\leq\overline{u}}S^{\prime\prime}(p)\cdot\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
ensuring that $I_{2,1}(\phi)$ belong to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega,\mathbb{R})$.
We continue by choosing $\alpha\in\mathopen{]}\frac{1}{2},1\mathclose{[}$ and then writing:
$|I_{2,2}(\phi)|$ $\leq\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|\mathinner{\mathrm{d}{x}}$
$\leq\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha-1}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\left(\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|\mathinner{\mathrm{d}{x}}\right)^{2}\right\}^{1/2}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\Delta x^{2}\right\}^{1/2}$
$\leq\sqrt{\frac{2TR}{\lambda}}\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha-1/2}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|^{2}\mathinner{\mathrm{d}{x}}\right\}^{1/2}$
[…]
We now take advantage of the compact embedding $\mathbf{C}^{0,\alpha}(\Omega,\mathbb{R})\subset\mathbf{W}^{1,p}(\Omega,\mathbb{R})$, $p=2/(1-\alpha)$ to be sure that $I_{2,2}(\phi)$ belongs to a compact subset of $\mathbf{W}^{-1,q_{3}}(\Omega,\mathbb{R})$, $q_{3}\coloneq 2/(1+\alpha)\in\mathopen{]}1,\frac{4}{3}\mathclose{[}$.
Finally, we estimate $I_{4}(\phi)$. Let $k\in[\underline{u},\overline{u}]$ and consider $(S_{i},Q_{i})_{i\in\{1,2\}}$ the entropy/entropy flux pairs defined in Lemma 3.11.
Because of the Rankine-Hugoniot condition,
$\left|[\![Q_{1}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|=|H(x_{j},k)-H(x_{j+1},k)|,$
from which we deduce
$|I_{4,1}(\phi)|\leq 2TR\;\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\cdot\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
ensuring that $I_{4,1}(\phi)$ belongs to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega,\mathbb{R})$. On the other hand,
$\left|[\![Q_{2}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|$ $=\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+\int_{k}^{\rho_{j+1/2}^{n,+}}\partial_{u}H(x_{j},\xi)^{2}-\partial_{u}H(x_{j+1},\xi)^{2}\mathinner{\mathrm{d}{\xi}}$
$\leq\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x$
$\leq L\underbrace{\int_{\rho^{-}}^{\rho^{+}}|\partial_{u}H(x_{j},\xi)|\mathinner{\mathrm{d}{\xi}}}_{Q_{2,1}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x,$
where $\rho^{-}\coloneq\min\{\rho_{j+1/2}^{n,-},\rho_{j+1/2}^{n,+}\}$ and $\rho^{+}\coloneq\max\{\rho_{j+1/2}^{n,-},\rho_{j+1/2}^{n,+}\}$. If $u\mapsto\partial_{u}H(x_{j},u)$ does not change sign in $\mathopen{]}\rho^{-},\rho^{+}\mathclose{[}$, then
$|Q_{2,1}|$ $=|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j},\rho_{j+1/2}^{n,+})|$
$\leq\underbrace{|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j+1},\rho_{j+1/2}^{n,+})|}_{=0}+|H(x_{j+1},\rho_{j+1/2}^{n,+})-H(x_{j},\rho_{j+1/2}^{n,+})|$
$\leq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\;\Delta x.$
Otherwise, assume to fix the ideas that $\rho^{-}\leq\alpha_{j}\leq\rho^{+}$. Then, setting $L_{H}\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{uu}H(x,p)|$ we write
$Q_{2,1}$ $=\int_{\rho^{-}}^{\rho^{+}}|\partial_{u}H(x_{j},\xi)-\partial_{u}H(x_{j},\alpha_{j})|\mathinner{\mathrm{d}{\xi}}$
$\leq L_{H}\int_{\rho^{-}}^{\rho^{+}}|\xi-\alpha_{j}|\mathinner{\mathrm{d}{\xi}}$
$=\frac{L_{H}}{2}\left((\rho^{+}-\alpha_{j})^{2}+(\rho^{-}-\alpha_{j})^{2}\right)$
$\leq L_{H}(\rho^{+}-\rho^{-})^{2}=L_{H}\;(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}.$
Taking advantage of Lemma 3.14, we conclude that
$\sum_{n=0}^{N}\Delta t\sum_{|j|\leq J}\left|[\![Q_{2}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|$
$\leq L\sum_{n=0}^{N}\Delta t\sum_{|j|\leq J}\left\{L_{H}\;(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\;\Delta x+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x\right\}$
$\leq L\left\{L_{H}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\right\}(\lambda c_{2}+2TR),$
which ensures that $I_{4,2}(\phi)$ belongs to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega)$.
To summarize, for any $i\in\{1,2\}$, $(\mu_{i})_{\Delta}$ belongs to a compact subset of $\mathbf{W}_{\mathbf{loc}}^{-1,q}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$, $q\coloneq\min\{q_{1},q_{2},q_{3}\}<2$. Additionally, $(\mu_{i})_{\Delta}$ is bounded in $\mathbf{W}_{\mathbf{loc}}^{-1,r}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for any $r\in\mathopen{]}2,+\infty\mathclose{[}$. Lemma 3.12 applies to ensure that for any $i\in\{1,2\}$, $(\mu_{i})_{\Delta}$ belongs to a compact subset of $\mathbf{H}_{\mathbf{loc}}^{-1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.   $\square$
Corollary 3.16.
Assume that $H$ satisfies (CH)–(CVX). Fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and let $(u_{\Delta})_{\Delta}$ be the sequence generated by the scheme described in Section 2.
Then, there exists a limit function $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ such that along a subsequence as $\Delta\to 0$, $(u_{\Delta})_{\Delta}$ converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for all $p\in[1,+\infty\mathclose{[}$ and a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to $u$.
Proof.  Follows from a combination of Lemma 3.13 and Lemma 3.15.   $\square$
3.3.2 Convergence
To prove that the limit function $u$ of Corollary 3.16 is the entropy solution, we will need the following technical result.
Lemma 3.17.
[…]
$\mathop{\rm sgn}(g_{\mathrm{\varepsilon}}-k){\ \underset{\mathrm{\varepsilon}\to 0}{\longrightarrow}\ }\mathop{\rm sgn}(g-k)\quad\text{a.e. in}\;\Omega.$
We can now prove our convergence statement.
Proof of (M2).  Let us split the proof into several steps.
Step 1.
Let us introduce the piecewise constant function: ${\rho_{\Delta}\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{[t^{n},t^{n+1}\mathclose{[}\times I_{j}}}$.
[…]
$|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|=|\mathcal{U}^{n}(t,x)-u_{j+1}^{n}|\leq|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}|,$
since $\mathcal{U}^{n}$ is the solution of a Riemann problem with left state $u_{j}^{n}$ and right state $u_{j+1}^{n}$ at $x=x_{j+1/2}$. Likewise, for all $(t,x)\in[t^{n},t^{n+1}\mathclose{[}\times\mathopen{]}x_{j},x_{j+1/2}\mathclose{[}$,
$|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|=|\mathcal{U}^{n}(t,x)-u_{j}^{n}|\leq|\rho_{j+1/2}^{n,-}-u_{j}^{n}|.$
Therefore, by Lemma 3.14, for all $N,K\in\mathbb{N}^{*}$,
$\sum_{n=0}^{N}\sum_{|j|\leq K}\int_{t^{n}}^{t^{n+1}}\int_{x_{j}}^{x_{j+1}}|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|^{2}\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$\leq\frac{1}{2}\sum_{n=0}^{N}\sum_{|j|\leq K}\left\{|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}|^{2}+|\rho_{j+1/2}^{n,-}-u_{j}^{n}|^{2}\right\}\Delta x\Delta t\leq\frac{c_{1}}{2}\Delta t,$
proving the claim. The claim ensures that $(\rho_{\Delta})_{\Delta}$ converges a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to $u$.
Step 2: Discrete entropy inequalities.
Let $k\in[\underline{u},\overline{u}]$. Under the CFL condition (2.3), we derive the following discrete entropy inequalities, consequence of the monotonicity of the scheme:
$\left(|u_{j}^{n+1}-k|-|u_{j}^{n}-k|\right)\Delta x+(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n})\Delta t$ (3.19)
$\leq-\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))\Delta t,$
where
$\Phi_{j+1/2}^{n}\coloneq F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\vee k,u_{j+1}^{n}\vee k)-F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\wedge k,u_{j+1}^{n}\wedge k).$
Indeed, on the one hand, by convexity,
$|u_{j}^{n+1}-H_{j}(k,k,k)|$ $\geq|u_{j}^{n+1}-k+\lambda(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j+1/2}(k,k))|$
$=|u_{j}^{n+1}-k|+\lambda\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k)).$
On the other hand, by monotonicity of the scheme,
$|u_{j}^{n+1}-H_{j}(k,k,k)|$ $\leq H_{j}(u_{j-1}^{n}\vee k,u_{j}^{n}\vee k,u_{j+1}^{n}\vee k)-H_{j}(u_{j-1}^{n}\wedge k,u_{j}^{n}\wedge k,u_{j+1}^{n}\wedge k)$
$=|u_{j}^{n}-k|-\lambda(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n}).$
Inequality (3.19) follows by combining these two estimates.
Step 3: Convergence.
Let $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ fix $T>0$, $R$ such that the support of $\varphi$ is included in $[0,T]\times[-R,R]$. Define
$\forall n\in\mathbb{N},\;\forall j\in\mathbb{Z},\quad\varphi_{j}^{n}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\varphi(t^{n},x)\mathinner{\mathrm{d}{x}}.$
With reference to Lemma 3.7, set $\mu\coloneq\sup_{k\in[\underline{u},\overline{u}]}\mu_{k}$ and fix $k\in[\underline{u},\overline{u}]$.
Multiply (3.19) by $\varphi_{j}^{n}$ and take the double sum for $n\in\mathbb{N}$ and $j\in\mathbb{Z}$. A summation by parts provides $A+B+C\geq 0$ with
$A$ :-$\coloneq\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}|u_{j}^{n}-k|(\varphi_{j}^{n}-\varphi_{j}^{n-1})\Delta x+\sum_{j\in\mathbb{Z}}|u_{j}^{o}-k|\varphi_{j}^{o}\Delta x:$
$B$ :-$\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\Phi_{j+1/2}^{n}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t:$
$C$ :-$\coloneq-\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))\varphi_{j}^{n}\Delta t.:$
Clearly,
$A{\ \underset{\Delta\to 0}{\longrightarrow}\ }\int_{0}^{+\infty}\int_{\mathbb{R}}|u-k|\partial_{t}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}u_{o}(x)\varphi(0,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
Consider now $B$. For all $n\in\mathbb{N}$ and $j\in\mathbb{Z}$, write
$\Phi_{j+1/2}^{n}$ $=\underbrace{F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\vee k,u_{j+1}^{n}\vee k)-F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\vee k,u_{j}^{n}\vee k)}_{B_{1}}$
$+\underbrace{F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\vee k,u_{j}^{n}\vee k)-H(x_{j+1},u_{j}^{n}\vee k)}_{B_{2}}+\underbrace{\Phi(x_{j+1},u_{j}^{n},k)-\Phi(x_{j},u_{j}^{n},k)}_{B_{3}}+\Phi(x_{j},u_{j}^{n},k)$
$+\underbrace{H(x_{j+1},u_{j}^{n}\wedge k)-F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\wedge k,u_{j}^{n}\wedge k)}_{B_{4}}+\underbrace{F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\wedge k,u_{j}^{n}\wedge k)-F_{\mathrm{int}}^{j+1/2}(u_{j}^{n}\wedge k,u_{j+1}^{n}\wedge k)}_{B_{5}}.$
We see that
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{1}+B_{5})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|$ $\leq 2L\biggl{(}T\|\partial_{xx}^{2}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}+\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{1}}\biggr{)}(\Delta x+\Delta t)$
$+2L\int_{0}^{+\infty}\int_{\mathbb{R}}|\rho_{\Delta}(t,x+\Delta x)-\rho_{\Delta}(t,x)|\cdot|\partial_{x}\varphi(t,x)|\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
Since $(\rho_{\Delta})_{\Delta}$ converges a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$, $(\rho_{\Delta}\partial_{x}\varphi)_{\Delta}$ is strongly compact in $\mathbf{L}^{1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. As a consequence of the Riesz-Fréchet-Kolmogorov compactness characterization, see [6, Chapter 4, Section 5],
[…]
On the other hand, using Lemma 3.7, we can estimate $B_{2}+B_{4}$ as:
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{2}+B_{4})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|\leq 2\mu T\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x.$
Regarding $B_{3}$, it is straightforward that
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}B_{3}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|\leq T\;\sup_{\underset{p,k\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}\Phi(x,p,k)|\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x.$
We now handle the last term in $B$. We have:
$\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\Phi(x_{j},u_{j}^{n},k)(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t$
$=\Delta t\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{I_{j}}\Phi(x_{j},u_{j}^{n},k)\left(\frac{\varphi(t^{n},x+\Delta x)-\varphi(t^{n},x)}{\Delta x}-\partial_{x}\varphi(t^{n},x)\right)\mathinner{\mathrm{d}{x}}$
$+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{t^{n}}^{t^{n+1}}\int_{I_{j}}\Phi(x_{j},u_{j}^{n},k)(\partial_{x}\varphi(t^{n},x)-\partial_{x}\varphi(t,x))\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{t^{n}}^{t^{n+1}}\int_{I_{j}}(\Phi(x_{j},u_{j}^{n},k)-\Phi(x,u_{j}^{n},k))\;\partial_{x}\varphi(t,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{0}^{+\infty}\int_{\mathbb{R}}\Phi(x,\rho_{\Delta},k)\partial_{x}\varphi(t,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$\leq T\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\Phi(x,p,k)|\cdot\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\Phi(x,p,k)|\cdot\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{1}}\Delta t$
$+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}\Phi(x,p,k)|\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{1}}\Delta x+\int_{0}^{+\infty}\int_{\mathbb{R}}\Phi(x,\rho_{\Delta},k)\partial_{x}\varphi(t,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},$
ensuring that
$B{\ \underset{\Delta\to 0}{\longrightarrow}\ }\int_{0}^{+\infty}\int_{\mathbb{R}}\Phi(x,u(t,x),k)\partial_{x}\varphi(t,x)\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
Finally, introduce $F_{\Delta}$ the continuous, linear affine function such that $F_{\Delta}(x_{j+1/2})=F_{\mathrm{int}}^{j+1/2}(k,k)$ for all $j\in\mathbb{Z}$. As a consequence of Lemma 3.7, $(F_{\Delta})_{\Delta}$ converges uniformly to $x\mapsto H(x,k)$ on $\mathbb{R}$. Moreover, since $(F_{\Delta})_{\Delta}$ is bounded in $\mathbf{Lip}(\mathbb{R},\mathbb{R})$, $(F_{\Delta}^{\prime})_{\Delta}$ converges weakly* in $\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, and the limit can only be $x\mapsto\partial_{x}H(x,k)$. We can now conclude with $C$. Write $C=C_{1}+C_{2}$, where
$C_{1}$ :-$\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))(\varphi_{j}^{n-1}-\varphi_{j}^{n})\Delta t:$
$C_{2}$ :-$\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{I_{j}}\mathop{\rm sgn}(u_{j}^{n}-k)F_{\Delta}^{\prime}(x)\varphi_{j}^{n}\Delta t.:$
On the one hand, using Lemma 3.7, we have
$|C_{1}|\leq\left(2\mu+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\left|\varphi_{j}^{n-1}-\varphi_{j}^{n}\right|\Delta x\Delta t\leq 2TR\left(2\mu+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\|\partial_{t}\varphi\|_{\mathbf{L}^{\infty}}\Delta t.$
On the other hand, we take advantage of Lemma 3.17 to obtain that
$\lim_{\Delta\to 0}C_{2}=-\int_{0}^{+\infty}\int_{\mathbb{R}}\mathop{\rm sgn}(u-k)\;\partial_{x}H(x,k)\varphi(t,x)\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},$
for any $k\in[\underline{u},\overline{u}]\backslash\mathcal{L}$. Then, repeat the argument from [25] to extend it for all $k\in[\underline{u},\overline{u}]$.
By passing to the limit $\Delta\to 0$ in $A+B+C\geq 0$, we proved that (1.1) holds, concluding the proof that $u$ is the entropy solution to (1).   $\square$
Acknowledgments.
The author thanks Nicolas Seguin for enlightening discussions.
The author acknowledges the PRIN 2022 project Modeling, Control and Games through Partial Differential Equations (D53D23005620006), funded by the European Union - Next Generation EU.